# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v2

@@ I Introduction
For decades, algorithms for finding near-optimal solution candidates to global optimization problems have been developed based on inspirations gleaned from nature. Famous examples include the use of biological evolution in the field of Evolutionary Computation [1, 2] (with examples such as Genetic Algorithms and Evolutionary Strategies), the swarming behavior of bird-like objects in Particle Swarm Optimization [3], or the foraging behavior of ants in Ant Colony Optimization [4]. For the highly advanced variants of such algorithms, impressive results have been reported for solving real-world optimization problems [5, 6, 7].
@@
Realizing that the laborious, expert driven approach for improving existing and inventing new algorithms has become quite inefficient, researchers have recently started to develop modular frameworks for arbitrarily combining components (modules) from algorithm classes into new variants - thereby creating combinatorial algorithm design spaces in which thousands to millions of algorithms can be generated, benchmarked, and optimized. Examples include the modular Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [11, 12], modular Differential Evolution [13], Particle Swarm Optimization [14], and a recent overview that summarizes all such approaches towards automated design [15].
@@
Although they often provide algorithm design spaces of millions of potential module combinations plus their hyperparameter search spaces, modular frameworks also need to be created by experts in the field, by carefully selecting the included modules and providing the full infrastructure for configuring and searching through such algorithm spaces. Moreover, they are limited to the design space created by the choices that have been made by the experts in selecting modules for inclusion.
@@
In this paper, we propose to overcome such limitations by using Large Language Models (LLMs) for automatically and iteratively evolving and optimizing the program code of such metaphor-based optimization algorithms. To test the approach, we use a specific optimization benchmarking tool to automatically evaluate the quality of the generated optimization algorithm and to provide a corresponding feedback to the LLM. The specific contributions of this work are as follows:
@@
  • We present the LLM-Evolutionary Algorithm (LLaMEA), an algorithm that uses an LLM to automatically generate and optimize high-quality metaphor-based optimization algorithms from scratch.
  • We couple LLaMEA with a benchmarking tool, IOHexperimenter [16], to automatically evaluate the quality of the generated optimization algorithm, for providing feedback to the LLM.
@@
  • We demonstrate that LLaMEA can generate new metaphor-based optimization algorithms that successfully compete with state-of-the-art algorithms on a standard set of benchmark test problems.
@@ II Related Work
The integration of Large Language Models (LLMs) into the optimization domain has recently received significant attention, resulting in various innovative methodologies.
Using optimization algorithms such as Genetic Algorithms for optimizing prompts (EvoPrompt) [17] has shown impressive results in outperforming human-engineered prompts. This idea of directly evolving the prompts by an Evolutionary Algorithm has also recently been used to demonstrate an automated engineering design optimization loop for finding 3D car shapes with optimal aerodynamic performance [18].
@@
In the Algorithm Evolution using Large Language Model (AEL) [22] approach, ideas from Evolutionary Algorithms are used explicitly to design novel optimization algorithms. AEL treats each algorithm as a solution candidate in a population, evolving them by asking the LLM to perform crossover and mutation operators on the algorithms generated by the LLM. This approach has shown superior performance over traditional heuristic methods and domain-specific models in solving small instances of the Traveling Salesperson Problem (TSP).
The direct application of LLMs as Evolution Strategies for optimization tasks represents another innovative approach, named EvoLLM [23]. In this proposal, the LLM generates the means of uni-variate normal distributions which are then used as sampling distributions for proposing solution vectors for black-box optimization tasks. To guide the LLM towards improvements, the best $m$ solution vectors from the best $k$ generations, each, are provided to the LLM, and it is prompted to generate the new mean values of the sampling distribution. A method such as EvoLLM leverages the LLM to perform the sampling based on the information summarized above, and it shows reasonable performance on a subset of eight of the BBOB set of benchmarking functions for continuous optimization [24] for one 5-dimensional instance of each problem. EvoLLM uses the idea of in-context learning [25, 26], which works by providing a set of examples and their corresponding scores to the LLM, which is then prompted to generate a score for a newly presented example.
@@
The recently proposed FunSearch approach (searching in the function space) [28] closes the loop and combines the LLM with a systematic evaluation of the quality of the generated programs, resulting in new solution algorithms for combinatorial optimization problems (cap-set problem, bin packing). In this approach, a distributed island-based Evolutionary Algorithm is used for maintaining diversity and prompt construction based on the already generated programs.
When it comes to leveraging LLMs for the design of metaheuristics, results from a manual prompting approach with GPT-4 for selecting, explaining and combining components of such algorithms have illustrated the capabilities of LLMs for generating such algorithms [29]. Automated generation and evaluation of their performance was not performed, however.
@@
In our approach, described in Section III, we propose to overcome such limitations by presenting a framework that can be used to generate any kind of algorithm as long there is a way to assess its quality automatically and it stays within the token limits of current LLMs. To achieve this goal, we combine in-context learning and an Evolutionary Algorithm-like iteration loop that allows the LLM to either refine (i.e., mutate) the best-so-far algorithm or redesign it completely. We combine this approach with a sound performance evaluation of the generated algorithms, based on the well-established IOHprofiler benchmark platform.
@@ III LLaMEA
The proposed LLaMEA framework (Figure 1) consists of four primary steps, which are iteratively repeated, following a similar structure as an Evolutionary Algorithm with one parent and one child (like a (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-EA) with an initialisation step and a general optimization loop of evaluation (in this case, by using IOHexperimenter [30]) and refinement or redesign of the algorithm, depending on the selection strategy (in case of a $(1+1)$-strategy, only improvements are accepted, while a $(1,1)$-strategy always accepts the newly generated algorithm).
@@
5:$(y_{t},\sigma_{t},e_{t})\leftarrow f(a_{t})$triangleright Evaluate mean fitness and std.-dev. of first program and catch errors if occurring
@@
8:     if $(1+1)$ then $\triangleright$ Construct new prompt, using best-so-far algorithm
@@
10:     $\triangleright$ Construct new prompt, using latest parent algorithm
@@
21:return $a_{b},y_{b}$triangleright Return best algorithm and its fitness
Our approach, which we abbreviate as (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-LLaMEA, is presented in more detail in Algorithm 1. Here, initialization is performed in lines 1-6 by prompting the LLM with a task description prompt $S$, evaluating the fitness $f(a_{t})$ of the generated program code (text) $a_{t}\in\mathcal{A}$, and memorizing the initial program $a_{t}$, its mean fitness $y_{t}$ and standard deviation of fitness $\sigma_{t}$ over multiple runs, and potentially some runtime error information $e_{t}$ as best solution $(a_{b},y_{b},\sigma_{b},e_{b})$ (see Section III-C for a variation of this approach, in which more detailed information is provided to LLaMEA).
The while-loop (line 7-20) iterates through $T$ (total budget) LLM calls to generate a new program each time (line 13). The prompt construction in lines 9 (in case of elitist selection) or line 11 (in case of non-elitist selection) is the essential step, generating a prompt $F$ that consist of a concatenation [^1] We use tuple notation here for clarity to show the components that compose the prompt string. of the following information: (i) Task description $S$; (ii) the algorithm names and mean fitness values of all previously generated algorithms; (iii) the current best algorithm’s $a_{b}$ (in case of elitist selection) or most recent algorithm’s $a_{t}$ (in case of non-elitist selection) complete information (code, mean fitness and standard deviation, execution error information) and (iv) a short task prompt $F_{0}$, with an instruction to the LLM. The new algorithm is generated by the LLM (line 13) and its fitness evaluated catching any runtime errors on the way (line 14), setting $y_{t+1}$ explicitly to zero (worst possible fitness) if runtime errors occurred (line 15). The best-so-far algorithm is updated if an improvement was found (line 17).
@@ III-A Starting Prompt
The initialisation of the optimization loop is crucial for the framework to work, as it sets the boundaries and rules that the LLM needs to operate with. Through experimentation, we found that including a small example code helps a lot in generating code without syntax and runtime errors. However, the example code could also bias the optimization search towards similar algorithms. In this work we therefore choose to use a simple implementation of random search as the example code to provide. The code of this random search algorithm is provided in our Zenodo repository [31]. The LLM receives an initial prompt with a specific set of criteria, domain expertise, and problem description. This starting point guides the LLM in generating an appropriate algorithm. The starting prompt has two main components, the specification of the role and general task, e.g., ”Your task is to design novel metaheuristic algorithms to solve black box optimization problems.”, and a detailed description of the problem to solve including a clear format of the expected response. Our detailed task prompt $S$ is given below:
  Detailed task prompt $S$     The optimization algorithm should handle a wide range of tasks,  which is evaluated on a large test suite of noiseless functions.  Your task is to write the optimization algorithm in Python code.  The code should contain one function `def __call__(self, f)`,  which should optimize the black box function `f` using  `budget` function evaluations.  The f() can only be called as many times as the budget allows.  An example of such code is as follows:  ```  <initial example code>  ```  Give a novel heuristic algorithm to solve  this task.  Give the response in the format:  # Name: <name of the algorithm>  # Code: <code>
@@ III-B Algorithm Synthesis (Initialization)
Using the prompt, the LLM generates the code for a new metaheuristic optimization algorithm, considering the constraints and guidance provided. The LLM should provide the answer in the format given, using MarkDown formatting for the code-block. The generated algorithm and name are extracted using regular expressions, including additional exception handling to capture small deviations from the asked format. In our experiments, 100% of the LLM responses lead to successfully extracted code. In addition, the LLM usually generates a small explanation in addition of what we ask, which we store for offline evaluation. In the case we cannot extract the code (which in practise never happened but can theoretically occur), we provide the LLM feedback that the response did not follow the provided format, trying to enforce the format for the next iteration.
@@
Once the algorithm is extracted we dynamically load the generated Python code and instantiate a version of the algorithm for evaluation. The loading and instantiating of the code can lead to syntax errors, which we capture and store to be provided in the refinement prompt. When these errors occur, the evaluation run cannot commence, and is therefore skipped. Evaluation metrics are set to the lowest possible value (00) for the particular candidate and error messages $e_{t}$ are stored for additional feedback to the LLM.
@@ III-C Evaluation
The generated algorithm is evaluated on the Black-Box Optimization Benchmark (BBOB) suite [24]. The suite consists of $24$ noiseless functions, with an instance generation mechanic to generate a diverse set of different optimization landscapes that share particular function characteristics. The suite is divided into $5$ function groups; separable functions, functions with low or moderate conditioning, high conditioning unimodal functions, multi-modal functions with strong global structure and multi-modal functions with weak global structure. The BBOB suite is considered a best-practice in the field of metaheuristic algorithm design as evaluation of newly proposed algorithms. For the evaluation feedback provided to the LLM, we summarize the performance on the whole BBOB suite by taking an average any-time performance metric, in our case the area over the convergence curve (AOCC) (see Equation 1). This results in one number $y_{t}$ representing the performance of the proposed algorithm on average over the complete benchmark suite, and its standard deviation, $\sigma_{t}$. In addition to aggregating these values over all functions, we also experiment with a more detailed feedback mechanism where the average AOCC and its standard deviation are returned for each function group, resulting in ten performance values $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. Theoretically, this approach should provide the LLM with information on what kind of functions the proposed solutions works well and on which it works less well.

### v3

@@ I Introduction
For decades, algorithms for finding near-optimal solution candidates to global optimization problems of the form
$\text{minimize}\;\mathcal{F}:\mathcal{S}\rightarrow\mathbb{R}$ (1)
where $\mathcal{S}\subseteq\mathbb{R}^{d}$ and $\mathcal{S}=\times_{i=1}^{d}[l_{i},u_{i}]$ is defined by box constraints ($l_{i}<u_{i}$, $\forall\,l_{i},u_{i}\in\mathbb{R}$), have been developed based on inspirations gleaned from nature. Famous examples include the use of biological evolution in the field of Evolutionary Computation [1, 2] (with examples such as Genetic Algorithms and Evolutionary Strategies), the swarming behavior of bird-like objects in Particle Swarm Optimization [3], or the foraging behavior of ants in Ant Colony Optimization [4]. For the highly advanced variants of such algorithms, impressive results have been reported for solving real-world optimization problems [5, 6, 7].
@@
Realizing that the laborious, expert driven approach for improving existing and inventing new algorithms has become quite inefficient, researchers have recently started to develop modular frameworks for arbitrarily combining components (modules) from algorithm classes into new variants - thereby creating combinatorial algorithm design spaces in which thousands to millions of algorithms can be generated, benchmarked, and optimized. Examples include the modular Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [11, 12], modular Differential Evolution [13], modular Particle Swarm Optimization [14], and a recent overview that summarizes all such approaches towards automated design [15].
@@
Although they often provide algorithm design spaces of millions of potential module combinations plus their hyperparameter search spaces, modular frameworks also need to be created by experts in the field, by carefully selecting the included modules and providing the full infrastructure for configuring and searching through such algorithm spaces. Moreover, they are limited to the design space created by the choices that have been made by the experts when selecting modules for inclusion.
@@
In this work, we propose to overcome such limitations by using Large Language Models (LLMs) within an evolutionary loop for automatically and iteratively evolving and optimizing the program code of such metaphor-based optimization algorithms for solving the optimization problem in Equation 1. To test the approach, we use a specific optimization benchmarking tool (IOHprofiler, consisting of the IOHexperimenter [16] module for systematically running algorithms on benchmark functions, and IOHanalyzer [17] for statistically analyzing the results of the runs) to automatically evaluate the quality of the generated optimization algorithm and to provide a corresponding feedback to the LLM. The specific contributions of this work are as follows:
@@
  • We present the LLM-Evolutionary Algorithm (LLaMEA), an evolutionary algorithm that uses an LLM to automatically generate and optimize high-quality metaphor-based optimization algorithms for solving the optimization problem as stated in Equation 1.
  • We couple LLaMEA with the benchmarking tool, IOHexperimenter [17], to automatically evaluate the quality of the generated optimization algorithm, for providing feedback to the LLM. IOHexperimenter is a well-established benchmarking framework that supports evaluating optimization heuristics for continuous optimization problems; see e.g. [18, 19, 20, 21, 22]. It allows for comparing a new optimization algorithm automatically against a well-known set of state-of-the-art algorithms on a wide set of benchmark functions in a statistically sound way. It is only this automated benchmarking approach that allows us to close the evolutionary loop that includes the LLM and to run it automatically.
@@
  • We demonstrate that LLaMEA can generate new metaphor-based optimization algorithms that successfully compete with or improve upon state-of-the-art algorithms on a standard set of 24 benchmark test problems (with multiple instances of each problem), the so-called black-box optimization benchmark BBOB [23], for the class of continuous optimization problem as stated in Equation 1. In particular, we show that the generated algorithms are competitive with the CMA-ES with default parameters, a hyperparameter-optimized CMA-ES, and Differential Evolution.
We also would like to emphasize that the goal of this research is not to generate a specific algorithm for combinatorial or constraint optimization problems, but rather to combine the use of modern, sound benchmarking techniques for measuring the performance of algorithms with the generative capabilities of LLaMEA for achieving reliable results of the runs for the continuous optimization task defined in Equation 1. Going beyond this research, future work will address other application domains, too.
@@ II Related Work
The integration of Large Language Models (LLMs) into the optimization domain has recently received significant attention, resulting in various innovative methodologies. We can distinguish three classes of solutions that integrate LLMs and Evolutionary Computation (EC); Prompt optimization by EC methods, LLMs as EC method and optimization of code generation.
Prompt optimization: Using optimization algorithms, such as Genetic Algorithms for optimizing prompts (EvoPrompt) [24] has shown impressive results in outperforming human-engineered prompts. This idea of directly evolving the prompts by an Evolutionary Algorithm has also recently been used to demonstrate an automated engineering design optimization loop for finding 3D car shapes with optimal aerodynamic performance [25].
@@
LLMs as EC: The direct application of LLMs as Evolution Strategies for optimization tasks represents another innovative approach, named EvoLLM [29]. In this proposal, the LLM generates the means of uni-variate normal distributions which are then used as sampling distributions for proposing solution vectors for black-box optimization tasks. To guide the LLM towards improvements, the best $m$ solution vectors from the best $k$ generations, each, are provided to the LLM, and it is prompted to generate the new mean values of the sampling distribution. A method such as EvoLLM leverages the LLM to perform the sampling based on the information summarized above, and it shows reasonable performance on a subset of eight of the BBOB set of benchmarking functions for continuous optimization [30] for one 5-dimensional instance of each problem. EvoLLM uses the idea of in-context learning [31, 32], which works by providing a set of examples and their corresponding scores to the LLM, which is then prompted to generate a score for a newly presented example.
Code generation: The recently proposed FunSearch approach (searching in the function space) [33] closes the loop and combines the LLM with a systematic evaluation of the quality of the generated programs, resulting in new solution algorithms for combinatorial optimization problems (cap-set problem, bin packing). In this approach, a distributed island-based Evolutionary Algorithm is used for maintaining diversity and prompt construction based on the already generated programs. In the approach called Algorithm Evolution using Large Language Model (AEL) [34] and also the Evolution of Heuristics (EoH) [35], ideas from Evolutionary Algorithms are used explicitly to design novel optimization heuristics. AEL and EoH treat each algorithm as a solution candidate in a population, evolving them by asking the LLM to perform crossover and mutation operators on the heuristics generated by the LLM. These approaches have shown superior performance over traditional heuristic methods and domain-specific models in solving small instances of the Traveling Salesperson Problem (TSP). The AEL approach, however, is very specific to the TSP and cannot be applied to the continuous optimization problem as in Equation 1, which is the subject of our work. The EoH approach can be generalized to the continuous optimization problem, and thus be compared with our proposed approach. Note that the EoH approach was designed to generate and optimize small pieces of code (single functions) and the black-box optimization algorithms we aim to generate can be complex, relatively large and consisting of a complete class with variables and functions.
@@
When it comes to leveraging LLMs for the design of black-box optimization metaheuristics, results from a manual prompting approach with GPT-4 for selecting, explaining and combining components of such algorithms have illustrated the capabilities of LLMs for generating such algorithms [37]. Automated generation and evaluation of their performance was not performed, however.
@@
In our approach, described in Section III, we propose to overcome such limitations by presenting a framework that can be used to generate any kind of algorithm as long there is a way to assess its quality automatically and it stays within the token limits of current LLMs. To achieve this goal, we combine in-context learning and an Evolutionary Algorithm-like iteration loop that allows the LLM to either refine (i.e., mutate) the best-so-far algorithm or redesign it completely. The proposed approach not only uses numerical feedback scores (as AEL and EoH do), but also debugging information or other textual feedback that can be automatically provided to steer the search towards better solutions. We combine this approach with a sound performance evaluation of the generated algorithms, based on the well-established IOHprofiler benchmark platform.
@@ III LLaMEA
The proposed Large Language Model Evolutionary Algorithm (LLaMEA) framework (Figure 1) consists of four primary steps, which are iteratively repeated, following a similar structure as an Evolutionary Algorithm with one parent and one child with an initialization step and a general optimization loop of evaluation (in this case, by using IOHexperimenter [38] to run the generated algorithms on a well defined set of benchmarking test functions) and refinement (mutation) or redesign of the algorithm. The selection strategy determines whether only improvements are accepted (in case of a $(1+1)$-strategy) or the newly generated algorithm is always accepted (in case of a $(1,1)$-strategy). Since both selection strategies are integrated into LLaMEA, we use the term (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-EA for describing both selection strategies in one notation.
@@
5:$(y_{t},\sigma_{t},e_{t})\leftarrow f(a_{t})$triangleright Evaluate mean quality and std.-dev. of first program and catch errors if occurring
@@
8:     if $(1+1)$ then $\triangleright$ $(1+1)$-Variant: Construct new prompt, using best-so-far algorithm
@@
10:     $\triangleright$ $(1,1)$-Variant: Construct new prompt, using latest parent algorithm
@@
21:return $a_{b},y_{b}$triangleright Return best algorithm and its quality
Our approach, which we therefore abbreviate as (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-LLaMEA, is presented in more detail in Algorithm 1. Here, initialization is performed in lines 1-6 by prompting the LLM with a task description prompt $S$ (see Section III-A for a description of $S$), evaluating the quality $f(a_{t})$ of the generated program code (text) $a_{t}\in\mathcal{A}$, and memorizing the initial program $a_{t}$, its mean quality $y_{t}$ and standard deviation of quality $\sigma_{t}$ over multiple runs, and potentially some runtime error information $e_{t}$ as best solution $(a_{b},y_{b},\sigma_{b},e_{b})$ (see Section III-C for a variation of this approach, in which more detailed information is provided to LLaMEA).
The while-loop (line 7-20) iterates through $T$ (total budget) LLM calls to generate a new program each time (line 13). The prompt construction in lines 9 (in case of elitist $(1+1)$-selection) or line 11 (in case of non-elitist $(1,1)$-selection) is the essential step, generating a prompt $F$ (see Section III-D for a description of the template of $F$) that consists of a concatenation [^1] We use tuple notation here for clarity to show the components that compose the prompt string. of the following information: (i) Task description $S$; (ii) the algorithm names and mean quality values of all previously generated algorithms; (iii) the current best algorithm’s $a_{b}$ (in case of $(1+1)$-selection) or most recent algorithm’s $a_{t}$ (in case of $(1,1)$-selection) complete information (code, mean quality and standard deviation, execution error information) and (iv) a short task prompt $F_{0}$, with an instruction to the LLM. The new algorithm is generated by the LLM (line 13) and its quality evaluated catching any runtime errors on the way (line 14), setting $y_{t+1}$ explicitly to zero (worst possible quality) if runtime errors occurred (line 15). The best-so-far algorithm is updated if an improvement was found (line 17).
It should be noted that in the general description of Algorithm 1 given above, we do not specify the quality measure $f:\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}$; $f(a)\rightarrow\max$ in detail, as LLaMEA is a generic concept. In Section III-C, we define the specific quality measure used here for generating metaheuristics.
@@ III-A Starting Prompt
The initialisation of the optimization loop is crucial for the framework to work, as it sets the boundaries and rules that the LLM needs to operate with. Through experimentation, we found that including a small example code into the task prompt $S$ helps a lot in generating code without syntax and runtime errors. However, the example code could also bias the search towards similar algorithms. In this work we therefore choose to use a simple implementation of random search as the example code to provide. The code of this random search algorithm is provided in our Zenodo repository [39]. The LLM receives an initial prompt with a specific set of criteria, domain expertise, and problem description. This starting point guides the LLM in generating an appropriate algorithm. The starting prompt provides the role description first, followed by a detailed description of the problem to solve including a clear format of the expected response. Our detailed task prompt $S$ is given below:
  Detailed task prompt $S$     Your task is to design novel metaheuristic algorithms to solve black box optimization problems.  The optimization algorithm should handle a wide range of tasks,  which is evaluated on a large test suite of noiseless functions.  Your task is to write the optimization algorithm in Python code.  The code should contain one function `def __call__(self, f)`,  which should optimize the black box function `f` using  `budget` function evaluations.  The f() can only be called as many times as the budget allows.  An example of such code is as follows:  ```  <initial example code>  ```  Give a novel heuristic algorithm to solve  this task.  Give the response in the format:  # Name: <name of the algorithm>  # Code: <code>
@@ III-B Algorithm Synthesis (Initialization)
Using the prompt $S$, the LLM generates the code for a new metaheuristic optimization algorithm, considering the constraints and guidance provided. The LLM should provide the answer in the format given, using MarkDown formatting for the code-block. The generated algorithm and its name are extracted using regular expressions, including additional exception handling to capture small deviations from the requested format. In our experiments, 100% of the LLM responses lead to successfully extracted code. In addition, the LLM usually generates a small explanation in addition to what we ask, which we store for offline evaluation. In the case that we cannot extract the code (which in practice never happened but can theoretically occur), we provide the LLM feedback that the response did not follow the provided format, trying to enforce the format for the next iteration.
@@
Once the algorithm is extracted we dynamically load the generated Python code and instantiate a version of the algorithm for evaluation. The loading and instantiating of the code can lead to syntax errors, which we capture and store to be provided in the refinement prompt $F$. When these errors occur, the evaluation run cannot commence, and is therefore skipped. Evaluation metrics are set to the lowest possible value (00) for the particular candidate and error messages $e_{t}$ are stored for additional feedback to the LLM.
@@ III-C Evaluation
The generated algorithm is evaluated on the Black-Box Optimization Benchmark (BBOB) suite (see [30] and the supplemental material for an overview of the functions). The suite consists of $24$ noiseless functions, with an instance generation mechanism to generate a diverse set of different optimization landscapes that share particular function characteristics. The suite is divided into five function groups: (i) separable functions, (ii) functions with low or moderate conditioning, (iii) high conditioning unimodal functions, (iv) multi-modal functions with strong global structure and (v) multi-modal functions with weak global structure. The BBOB suite is considered a best-practice in the field of metaheuristic algorithm design for the evaluation of newly proposed algorithms. For the evaluation feedback provided to the LLM, we summarize the performance on the whole BBOB suite by taking an average any-time performance metric, namely the normalized Area Over the Convergence Curve (AOCC) (see Equation 2). This results in one number $y_{t}$ representing aggregated performance of the proposed algorithm over the complete benchmark suite, including multiple instances per function, and its standard deviation $\sigma_{t}$ over multiple runs of the algorithm.
Each generated algorithm is run for a fixed budget $B$ of function evaluations, and the IOHexperimenter suite makes sure that the algorithm is terminated after using the full budget of evaluations. In practice, this mechanism ensures that generated algorithms do not exceed the budget and always terminate. In addition to aggregating these values over all functions, we also experiment with a more detailed feedback mechanism where the average AOCC and its standard deviation are returned for each function group, resulting in ten performance values $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$ (two per function group). Theoretically, this approach should provide the LLM with information on what kind of functions the proposed algorithm works well and on which it works less well.